# UDPNet: Unleashing Depth-based Priors for Robust Image Dehazing

### v1

@@ Abstract
Image dehazing has witnessed significant advancements with the development of deep learning models. However, a few methods predominantly focus on single-modal RGB features, neglecting the inherent correlation between scene depth and haze distribution. Even those that jointly optimize depth estimation and image dehazing often suffer from suboptimal performance due to inadequate utilization of accurate depth information. In this paper, we present UDPNet, a general framework that leverages depth-based priors from large-scale pretrained depth estimation model DepthAnything V2 to boost existing image dehazing models. Specifically, our architecture comprises two typical components: the Depth-Guided Attention Module (DGAM) adaptively modulates features via lightweight depth-guided channel attention, and the Depth Prior Fusion Module (DPFM) enables hierarchical fusion of multi-scale depth map features by dual sliding-window multi-head cross-attention mechanism. These modules ensure both computational efficiency and effective integration of depth priors. Moreover, the intrinsic robustness of depth priors empowers the network to dynamically adapt to varying haze densities, illumination conditions, and domain gaps across synthetic and real-world data. Extensive experimental results demonstrate the effectiveness of our UDPNet, outperforming the state-of-the-art methods on popular dehazing datasets, such as 0.85 dB PSNR improvement on the SOTS dataset, 1.19 dB on the Haze4K dataset and 1.79 dB PSNR on the NHR dataset. Our proposed solution establishes a new benchmark for depth-aware dehazing across various scenarios. Pretrained models and codes will be released at our project https://github.com/Harbinzzy/UDPNet.
journal: International Journal of Computer Vision
1 Introduction
Figure 1: Illustration of Representative Image Dehazing Frameworks. (a) Approaches based on deep learning Dong et al (2020); Hong et al (2020); Wu et al (2021); Cui et al (2023b, 2024a); Liu et al (2021b); Zamir et al (2022); Guo et al (2022); Qiu et al (2023); Jin et al (2025) or physical models He et al (2010); Zhu et al (2015); Berman et al (2018); He et al (2012); Ju et al (2019); Zhao (2021) fail in complex distortions due to idealized assumptions and limited data, resulting in unsatisfactory dehazing outcomes and noticeable artifacts. (b) The collaborative mutual promotion between depth estimation and image dehazing Zhang et al (2024) requires alternate training of two networks, and has lower performance in outdoor and real-world scenarios. (c) The depth-consistent dehazing network Wang et al (2024a) according to the differences fails to make good use of depth information. Although the perception quality has been enhanced, the distortion metric is poor. Both (b) and (c) suffer from the inevitable error in depth estimation. (d) Our UDPNet utilizes the robust and reliable depth estimation model trained on large-scale datasets and effectively integrates it into a multi-scale hierarchical network, achieving excellent generalization.
Image dehazing, a fundamental low-level image restoration task, plays an essential preprocessing role in enhancing the reliability of downstream high-level vision systems (e.g., object detection, semantic segmentation). According to the ASM Middleton (1957), the degradation process of hazy images can be mathematically described as:
@@ 1 Introduction
where $(x,y)$ represents the pixel coordinate, $c\in\{r,g,b\}$ is the color channel, $H$ denotes the hazy image, $J$ denotes the corresponding clear image, $A$ is the atmospheric light. The transmission map is usually modeled as $t(x,y)=e^{-\beta\cdot d(x,y)}$ with the scattering coefficient $\beta$ and the scene depth $d(x,y)$. It follows from the above equation that the estimation accuracy of the transmission map and atmospheric light is crucial to traditional dehazing methods Zhu et al (2015); Berman et al (2018); He et al (2012); Ju et al (2019); Zhao (2021), which rely on reasonable assumptions or statistical priors (see Fig. 1 (a)). However, the complexity of real-world environments often leads to violations of the underlying priors, thereby reducing the robustness of traditional dehazing methods. For instance, He et al. He et al (2010) assumed that at least one color channel has very low intensity in most haze-free patches, which may not hold in bright regions such as the sky or objects with high reflectance, resulting in color distortions. Similarly, priors based on color-line distributions Zhu et al (2015) or contrast assumptions Ju et al (2019) can break down in cases where the scene contains complex textures, non-uniform illumination, or artificial light sources. Moreover, the estimation of atmospheric light can be unreliable in scenes with strong sunlight, multiple scattering, or dense haze, further amplifying the reconstruction error. These limitations highlight the inherent fragility of prior-based dehazing when applied to real-world conditions.
@@
With the advent of deep learning, various variations of Convolutional Neural Networks (CNNs) Dong et al (2020); Hong et al (2020); Wu et al (2021); Cui et al (2023b, 2024a) and Transformers Liu et al (2021b); Zamir et al (2022); Guo et al (2022); Qiu et al (2023); Jin et al (2025) have emerged for single image dehazing by learning the strong statistical regularities. They typically compute a sequence of features from the single-modal RGB input images and directly reconstruct the clear ones (see Fig. 1 (a)). Such approaches have achieved state-of-the-art results on benchmark datasets by leveraging advanced architectural designs such as multi-scale information fusion Ren et al (2020); Liu et al (2019); Cui et al (2023b, 2024a), multi-stage pipelines, sophisticated variants of convolution Liu et al (2020); Luo et al (2023); Cui et al (2024b), attention mechanisms Liu et al (2019); Cui et al (2024a). Nevertheless, dehazing remains an ill‑posed inverse problem in which feature estimation errors can markedly degrade restoration quality. Since RGB observations alone provide incomplete and often ambiguous cues under heavy haze, relying exclusively on them limits the model’s ability to resolve uncertainty and correct erroneous predictions. Inspired by the human visual system’s capacity to integrate and reconcile multisensory information for robust perception, we argue that effective dehazing requires the joint exploitation of complementary modalities (e.g., depth maps, FIR Du et al (2025), NIR Li et al (2025); Feng et al (2013), SAR Zhao et al (2025), semantic maps Zhang et al (2021); Jin et al (2023b)) to enhance robustness and generalization in haze removal. The Atmospheric Scattering Model (ASM) Middleton (1957) further provides a theoretical basis, as haze density is intrinsically correlated with scene depth, suggesting a natural synergy between dehazing and depth estimation. Leveraging depth-based priors can thus supply structural guidance and holistic awareness of haze distribution. Following this insight, several recent works Wang et al (2024a); Zhang et al (2024) integrated depth into dehazing frameworks. For instance, DCMPNet Zhang et al (2024) (see Fig. 1 (b)) incorporated the depth estimation and image dehazing model within a unified framework in a dual-task-driven manner. SelfPromer Wang et al (2024a) (see Fig. 1 (c)) extracted the depth difference features between hazy residuals and clear counterparts as prompts to improve haze removal.
Figure 2: Previous image dehazing methods often lack robustness to changes in scene conditions. In contrast, our general model effectively restores hazy images under varying illumination, varying scenes, haze distributions, and haze densities. Zoom in for more details.
@@
However, there still exist three main issues in practical image dehazing: (1) The acquired depth information is usually inaccurate. Prior studies have shown that larger datasets and more powerful models provide richer cues beyond basic labeling. Yet existing depth-based methods, whether relying on pretrained depth estimation models Birkl et al (2023) with limited training data or lightweight models trained from scratch Yang and Zhang (2022); Zhang et al (2024), still struggle to cope with complex hazy scenes such as nonhomogeneous, dense, or nighttime haze. (2) The potential of depth information has not been fully exploited. Although the difference self-prompt approach Wang et al (2024a) improved perceptual quality, it still performed poorly on distortion-oriented metrics. Similarly, directly integrating depth features at the tail of a UNet backbone Zhang et al (2024) failed to yield satisfactory results. More recently, PromptHaze Ye et al (2025) adopted a relatively straightforward design by injecting depth prompt features as an additional input, which limited its ability to capture the complex interactions between depth and haze. (3) A lack of universal high-performing dehazing models for both daytime and nighttime dehazing. Real-world nighttime haze is further complicated by multiple scattering, uneven illumination, glow, blur, and hidden noise, which make it more challenging than daytime haze. Existing state-of-the-art daytime dehazing methods Liu et al (2021b); Zamir et al (2022); Guo et al (2022); Qiu et al (2023); Jin et al (2025) and nighttime counterparts Jin et al (2023a); Liu et al (2023) are typically developed in isolation, and none can effectively handle both scenarios in a single framework. Thus, developing a universal model that can jointly address daytime haze, nighttime haze, and diverse haze conditions is still an urgent yet unresolved challenge, as illustrated in Fig. 2.
@@
To address these challenges, we turn to the fundamental limitations posed by depth estimation and its integration with visual features in the context of image dehazing. Depth estimation, while a promising cue, is often hindered by inherent inaccuracies and limited generalization across complex hazy scenes (Issue 1). In response, we employ DepthAnything V2 Yang et al (2024b), a state-of-the-art pretrained model, which offers robust and generalizable depth cues. To fully exploit the potential of depth as a structural prior (Issue 2), we propose two novel modules that effectively integrate depth with visual representations. The Depth-Guided Attention Module (DGAM) adaptively modulates feature maps via depth-informed channel attention, ensuring efficient interaction between depth and image content. Complementing this, the Depth Prior Fusion Module (DPFM) enables hierarchical fusion of multi-scale depth features with visual representations, facilitating richer structural guidance and enhancing robustness in dehazing tasks. This dual-module approach bridges the gap between depth and image representations, providing a unified solution for diverse hazy scenarios.
@@ 1. $\bullet$
    We introduce the large-scale pretrained depth estimation model for image dehazing. The incorporation of powerful priors enables our network to generalize to a wide range of hazy image scenarios.
@@ 3. $\bullet$
    Our model, UDPNet, achieves state-of-the-art performance on popular dehazing datasets, exhibiting strong robustness across both daytime/nighttime conditions and synthetic/real-world scenarios, while also demonstrating its potential in other image restoration tasks.
@@
The remainder of this paper is organized as follows. Section 2 reviews existing daytime and nighttime image dehazing methods, with a particular focus on depth-based approaches. Section 3 presents our observations and motivation, along with preliminary experiments for validation. Section 4 details the proposed UDPNet model, and Section 5 presents experimental results across various datasets along with discussions on limitations. Finally, Section 6 concludes the paper.
2 Related Work
2.1 Image Dehazing
@@
Image dehazing methods are typically divided into prior-based, learning-based, and hybrid approaches. Because of its ill-posed nature, prior-based methods rely on the atmospheric scattering model and handcrafted priors, such as the dark channel prior (DCP) He et al (2010), color attenuation prior (CAP) Zhu et al (2015), and BCCR Meng et al (2013), to estimate the transmission map and atmospheric light. While effective under ideal conditions that meet their assumptions, these approaches struggle to generalize in complex haze scenarios. In contrast, learning-based methods leverage data-driven architectures such as CNNs Krizhevsky et al (2012) and Transformers Vaswani et al (2017) to achieve superior dehazing performance. Some early models focus on learning transmission maps Ren et al (2016); Cai et al (2016) or jointly learning transmission maps and atmospheric light Pang et al (2018); Zhang and Patel (2018); Li et al (2021). In parallel, RDN Li et al (2020) and Dehaze-RetinexGAN Wang et al (2025b) embed the Retinex-based decomposition model into the dehazing network to improve generalization ability.
@@
However, these methods are susceptible to cumulative errors arising from inaccurate estimations. To avoid this, more recent works Liu et al (2019); Wang et al (2021); Chen et al (2024); Cui et al (2024b) incorporate hierarchical multi-scale representations, attention mechanisms, large-kernel convolutions, etc, enabling effective dehazing without reliance on the physical model. GridDehazeNet Liu et al (2019) employs a three-stage attention-based multi-scale framework to achieve end-to-end dehazing. EPDN Qu et al (2019), a pioneering framework embedding GAN and the enhancer module, generates perceptually pleasing images with realistic colors and fine details. FFA-Net Qin et al (2020) introduces the feature attention (FA) and attention-based feature fusion (FFA) to effectively integrate information across different levels. Cui et al. Cui et al (2023b, 2024a); Su et al (2025) design the hierarchical multi-input multi-output pure convolutional network to recover image sharpness and textures. C2PNet Zheng et al (2023) and PHATNet Tsai et al (2025) incorporate physics priors into the feature space, enhancing interpretability by aligning feature representations with the hazing process. Dehamer Guo et al (2022), DehazeFormer Song et al (2023) and MB-TaylorFormer Qiu et al (2023) leverage Transformers to model long-range dependencies. Concurrently, with the advancement of data-driven methods in dehazing performance, network complexity has also increased. To mitigate the reliance on large-scale training data, hybrid approaches Mo et al (2022) have been proposed, which integrate inherent image priors with the representation power of deep neural networks. For example, PGH2Net Su et al (2025) introduces triple priors (i.e., bright channel prior, dark channel prior and histogram equalization prior) to UNet-like architecture for efficient image dehazing. Methods like Dudhane and Murala (2019); Fang et al (2025) utilize transformation of color space (YCbCr) to guide the RGB space. Furthermore, SelfPromer Wang et al (2024a) and PromptHaze Ye et al (2025) explore the fruitful combination of depth-based priors with deep dehazing network.
@@
In addition to these architectural advances, several complementary directions have emerged, including methods for nighttime dehazing Yan et al (2020); Jin et al (2023a), haze image generation approaches Wang et al (2025a); Wu et al (2023) for real-world scenarios, and recent efforts toward universal image restoration Jiang et al (2025). For example, NightHazeFormer Liu et al (2023) integrates dual priors into the transformer decoder, effectively guiding the network to learn abundant prior features. NightHaze Lin et al (2025) introduces a self-prior learning paradigm, leveraging MAE-like representations to improve nighttime dehazing performance. Wang et al. Wang et al (2025a) propose HazeGen and DiffDehaze, a diffusion-based hazing–dehazing pipeline. Together, these efforts broaden the scope of dehazing research and lay a solid foundation for addressing more challenging restoration tasks.
2.2 Depth Priors for Image Dehazing

### v2

@@ Abstract
Image dehazing has witnessed significant advancements with the development of deep learning models. However, most existing methods focus solely on single-modal RGB features, neglecting the inherent correlation between scene depth and haze distribution. Even those that jointly optimize depth estimation and image dehazing often suffer from suboptimal performance due to inadequate utilization of accurate depth information. In this paper, we present UDPNet, a general framework that leverages depth-based priors from a large-scale pretrained depth estimation model DepthAnything V2 to boost existing image dehazing models. Specifically, our architecture comprises two key components: the Depth-Guided Attention Module (DGAM) adaptively modulates features via lightweight depth-guided channel attention, and the Depth Prior Fusion Module (DPFM) enables hierarchical fusion of multi-scale depth map features by dual sliding-window multi-head cross-attention mechanism. These modules ensure both computational efficiency and effective integration of depth priors. Moreover, the depth priors empower the network to dynamically adapt to varying haze densities, illumination conditions, and domain gaps across synthetic and real-world data. Extensive experimental results demonstrate the effectiveness of our UDPNet, outperforming the state-of-the-art methods on popular dehazing datasets, with PSNR improvements of 0.85 dB on SOTS-indoor, 1.19 dB on Haze4K, and 1.79 dB on NHR. Our proposed solution establishes a new benchmark for depth-aware dehazing across various scenarios. Pretrained models and codes are released at our project  https://github.com/Harbinzzy/UDPNet.
I Introduction
Image dehazing, a fundamental low-level image restoration task, plays an essential preprocessing role in enhancing the reliability of downstream high-level vision systems (e.g., object detection, semantic segmentation). According to the Atmospheric Scattering Model (ASM) [middleton1957vision], the degradation process of hazy images can be mathematically described as:
@@ I Introduction
where $(x,y)$ represents the pixel coordinate, $c\in\{r,g,b\}$ denotes the color channel, $H$ denotes the hazy image, $J$ denotes the corresponding clear image, $A$ is the atmospheric light. The transmission map is usually modeled as $t(x,y)=e^{-\beta\cdot d(x,y)}$ with the scattering coefficient $\beta$ and the scene depth $d(x,y)$. It follows from the above equation that the estimation accuracy of the transmission map and atmospheric light is crucial to traditional dehazing methods [zhu2015fast, he2012guided, ju2019idgcp, he2010single], which rely on reasonable assumptions or statistical priors (see Fig. 1 (a)). However, the complexity of real-world environments often leads to violations of the underlying priors, thereby reducing the robustness of traditional dehazing methods. For instance, He et al. [he2010single] assumed that at least one color channel has very low intensity in most haze-free patches, which may not hold in bright regions such as the sky or objects with high reflectance, resulting in color distortions. Similarly, priors based on color-line distributions [zhu2015fast] or contrast assumptions [ju2019idgcp] can break down in cases where the scene contains complex textures, non-uniform illumination, or artificial light sources. Moreover, the estimation of atmospheric light can be unreliable in scenes with strong sunlight, multiple scattering, or dense haze, further amplifying the reconstruction error. These limitations highlight the inherent fragility of prior-based dehazing when applied to real-world conditions.
Figure 1: Illustration of Representative Image Dehazing Frameworks. (a) Approaches based on deep learning [dong2020multi, wu2021contrastive, cui2024revitalizing, qiu2023mb, jin2025mb] or physical models [he2010single, zhu2015fast, he2012guided, ju2019idgcp]. (b) The collaborative mutual promotion between depth estimation and image dehazing [zhang2024depth]. (c) The depth-consistent prompt dehazing network [wang2024selfpromer]. (d) Our UDPNet utilizes the robust and reliable depth estimation model trained on large-scale datasets and effectively integrates it into a multi-scale hierarchical network, achieving excellent generalization.
@@
With the advent of deep learning, various variations of Convolutional Neural Networks (CNNs) [dong2020multi, wu2021contrastive, cui2023image, cui2024revitalizing] and Transformers [liu2021swin, zamir2022restormer, guo2022image, qiu2023mb, jin2025mb] have emerged for image dehazing by learning the strong statistical regularities. They typically compute a sequence of features from the single-modal RGB inputs and directly reconstruct the clear ones (see Fig. 1 (a)). Such approaches have achieved state-of-the-art results on benchmark datasets by leveraging advanced architectural designs such as multi-scale information fusion [liu2019griddehazenet, cui2023image, cui2024revitalizing], multi-stage pipelines, sophisticated variants of convolution [liu2020trident, luo2023lkd, cui2024omni], attention mechanisms [liu2019griddehazenet, cui2024revitalizing]. Nevertheless, dehazing remains an ill‑posed inverse problem in which feature estimation errors can markedly degrade restoration quality. Since RGB observations alone provide incomplete and often ambiguous cues under heavy haze, relying exclusively on them limits the model’s ability to resolve uncertainty and correct erroneous predictions. Inspired by the human visual system’s capacity to integrate and reconcile multisensory information for robust perception, we argue that effective dehazing requires the joint exploitation of complementary modalities (e.g., depth maps [zhang2024depth], near-infrared [li2025hierarchical] or semantic maps [zhang2021semantic]) to enhance robustness and generalization in haze removal. The ASM further provides a theoretical basis, as haze density is intrinsically correlated with scene depth, suggesting a natural synergy between dehazing and depth estimation. Leveraging depth-based priors can thus supply structural guidance and holistic awareness of haze distribution. Following this insight, several recent works [wang2024selfpromer, zhang2024depth] integrated depth into dehazing frameworks. For instance, DCMPNet [zhang2024depth] (see Fig. 1 (b)) incorporated the depth estimation and image dehazing model within a unified framework in a dual-task-driven manner, while exhibiting reduced effectiveness in outdoor and real-world scenarios. SelfPromer [wang2024selfpromer] (see Fig. 1 (c)) extracted the depth difference features between hazy and clear images as prompts to improve haze removal. Although the perception quality has been enhanced, the distortion metric is poor.
Figure 2: Comparisons of dehazing methods across varying conditions. Our framework demonstrates consistent performance improvements compared to representative methods.
@@
However, there still exist three main challenges in practical image dehazing: (1) The acquired depth information is usually inaccurate. Prior studies have shown that larger datasets and more powerful models provide richer cues beyond basic labeling. Yet existing depth-based methods, whether relying on pretrained depth estimation models [birkl2023midas] with limited training data or lightweight models trained from scratch [yang2022depth, zhang2024depth], still struggle to cope with complex hazy scenes such as nonhomogeneous, dense, or nighttime haze. (2) The potential of depth information has not been fully exploited. Although the difference self-prompt approach [wang2024selfpromer] improved perceptual quality, it still performed poorly on distortion-oriented metrics. Similarly, directly integrating depth features at the tail of a UNet backbone [zhang2024depth] failed to yield satisfactory results. More recently, PromptHaze [ye2025prompthaze] adopted a relatively straightforward design by injecting depth prompt features as an additional input, which limited its ability to capture the complex interactions between depth and haze. (3) Consistently high-performing dehazing across both daytime and nighttime remains challenging. Real-world nighttime haze is further complicated by multiple scattering, uneven illumination, glow, blur, and hidden noise, which make it more challenging than daytime haze. Existing state-of-the-art daytime dehazing methods [liu2021swin, zamir2022restormer, guo2022image, qiu2023mb] and nighttime counterparts [jin2023enhancing, liu2023nighthazeformer] are not explicitly designed to handle both daytime and nighttime scenarios. Thus, developing a universal model that can jointly address daytime haze, nighttime haze, and diverse haze conditions is still an urgent challenge, as illustrated in Fig. 2.
@@
To address these challenges, we turn to the fundamental limitations posed by depth estimation and its integration with visual features in the context of image dehazing. Depth estimation, while a promising cue, is often hindered by inherent inaccuracies and limited generalization across complex hazy scenes (Issue 1). In response, we employ DepthAnything V2 [depthanything_v2], a state-of-the-art pretrained model, which offers robust and generalizable depth cues. To fully exploit the potential of depth as a structural prior (Issue 2), we propose two lightweight plug-and-play modules that effectively integrate depth with visual representations. The Depth-Guided Attention Module (DGAM) adaptively modulates feature maps via depth-informed channel attention, ensuring efficient interaction between depth and image content. Complementing this, the Depth Prior Fusion Module (DPFM) enables hierarchical fusion of multi-scale depth features with visual representations, facilitating richer structural guidance and enhancing robustness in dehazing tasks. This dual-module approach bridges the gap between depth and image representations, providing a unified solution for diverse hazy scenarios.
@@ 1. $\bullet$
    We introduce the large-scale pretrained depth estimation model for image dehazing. The systematical incorporation of priors enables our network to generalize to a wide range of hazy image scenarios.
@@ 3. $\bullet$
    Our UDPNet model achieve state-of-the-art performance on various dehazing datasets, exhibiting consistent improvements across both daytime/nighttime conditions and synthetic/real-world scenarios, while also demonstrating potential in other image restoration tasks.
@@
The remainder of this paper is organized as follows. Sec. II reviews existing image dehazing methods, with a particular focus on depth-based approaches. Sec. III presents our observations and motivation, along with preliminary experiments. Sec. IV details the proposed UDPNet, and Sec. V presents experimental results across popular datasets along with discussions on limitations. Finally, Sec. VI concludes the paper.
II Related Work
II-A Image Dehazing
@@
Image dehazing methods are typically divided into prior-based, learning-based, and hybrid approaches. Because of its ill-posed nature, prior-based methods rely on the atmospheric scattering model and handcrafted priors, such as the dark channel prior (DCP) [he2010single], color attenuation prior (CAP) [zhu2015fast], and BCCR [meng2013efficient], to estimate the transmission map and atmospheric light. While effective under ideal conditions that meet their assumptions, these approaches struggle to generalize in complex haze scenarios. In contrast, learning-based methods leverage data-driven architectures such as CNNs and Transformers to achieve superior dehazing performance. Some early models focus on learning transmission maps [ren2016single, cai2016dehazenet] or jointly learning transmission maps and atmospheric light [pang2018visual, zhang2018densely]. In parallel, Dehaze-RetinexGAN [wang2025dehaze] embed the Retinex-based decomposition model into the dehazing network to improve generalization ability.
@@
However, these methods are susceptible to cumulative errors arising from inaccurate estimations. To avoid this, more recent works [liu2019griddehazenet, wang2021ensemble, chen2024dea, cui2024omni] incorporate hierarchical multi-scale representations, attention mechanisms, large-kernel convolutions, enabling effective dehazing without reliance on the physical model. GridDehazeNet [liu2019griddehazenet] employs a three-stage attention-based multi-scale framework to achieve end-to-end dehazing. EPDN [qu2019enhanced], a pioneering framework embedding GAN and the enhancer module, generates perceptually pleasing images with realistic colors and fine details. FFA-Net [qin2020ffa] introduces the feature attention (FA) and attention-based feature fusion (FFA) to effectively integrate information across different levels. Cui et al. [cui2023image, cui2024revitalizing, su2025prior] design the hierarchical multi-input multi-output pure convolutional network to recover image sharpness and textures. C2PNet [zheng2023curricular] and PHATNet [tsai2025phatnet] incorporate physics priors into the feature space, enhancing interpretability by aligning feature representations with the hazing process. Dehamer [guo2022image], DehazeFormer [song2023vision] and MB-TaylorFormer [qiu2023mb] leverage Transformers to model long-range dependencies. Concurrently, with the advancement of data-driven methods in dehazing performance, network complexity also increases. To mitigate the reliance on large-scale training data, hybrid approaches [mo2022dca] were proposed, integrating inherent image priors with the representation power of deep neural networks. For example, PGH2Net [su2025prior] introduced triple priors (i.e., bright channel prior, dark channel prior and histogram equalization prior) to UNet-like architecture. Methods like  [fang2025guided] utilize transformation of color space (YCbCr) to guide the RGB space. Furthermore, SelfPromer [wang2024selfpromer] and PromptHaze [ye2025prompthaze] explore the fruitful combination of depth-based priors with deep dehazing network.
@@
In addition to these architectural advances, several complementary directions have emerged, including methods for nighttime dehazing [yan2020nighttime, jin2023enhancing], haze image generation approaches [wang2025learning, wu2023ridcp] for real-world scenarios, and recent efforts toward universal image restoration [jiang2024survey]. For example, NightHazeFormer [liu2023nighthazeformer] integrates dual priors into the transformer decoder, effectively guiding the network to learn abundant prior features. NightHaze [lin2025nighthaze] introduces a self-prior learning paradigm, leveraging MAE-like representations to improve nighttime dehazing performance. Wang et al. [wang2025learning] propose HazeGen and DiffDehaze, a diffusion-based hazing–dehazing pipeline. Together, these efforts broaden the scope of dehazing research and lay a solid foundation for addressing more challenging restoration tasks.
II-B Depth Priors for Image Dehazing